OpenStudy (anonymous):

Evaluate: 4 − 35 / 7 + 32

OpenStudy (anonymous):

= 4 - 5 + 32 = (-1) + 32 = 31

OpenStudy (anonymous):

the 2 on the 32 is sapost to be above the 3

OpenStudy (anonymous):

 like that

OpenStudy (anonymous):

3^2?

OpenStudy (anonymous):

PEMDAS 4 - 35/7 + 9 4 - 5 + 9 4 - 14 -10

OpenStudy (anonymous):

= 4 − 35 / 7 + 3^2 = 4 - 5 + 9 = (-1) + 9 = 8

OpenStudy (anonymous):

I assume it is 3 to the power of 2.

OpenStudy (anonymous):

A comes before S... PEMDAS

OpenStudy (anonymous):

is it -10 or 8 ??

OpenStudy (anonymous):

It's -10

OpenStudy (anonymous):

thankss

OpenStudy (anonymous):

Sure :-)

OpenStudy (anonymous):

Yes either way its (-5) + 9...then + 4 = 8

OpenStudy (anonymous):

-10 is not right.

OpenStudy (anonymous):

lol ok.